OpenStudy (anonymous):

Doug is a salesperson in a retail store and earns $75 per week, plus 15% of his weekly sales. If Doug makes $555 one week, what are his sales for that week? $3,200 $627 $566 $1,914 Thanks for all the help :D

OpenStudy (anonymous):

The answer is $3200. =D

OpenStudy (anonymous):

:o how =3

OpenStudy (anonymous):

75 + 15x/100 = 555 15x/100 = 555-75 x = 480 * 100 / 15 Which is 3200.

OpenStudy (anonymous):

You got it? =)